# Few-shot Online Anomaly Detection and Segmentation

### v1

@@ Abstract
Detecting anomaly patterns from images is a crucial artificial intelligence technique in industrial applications. Recent research in this domain has emphasized the necessity of a large volume of training data, overlooking the practical scenario where, post-deployment of the model, unlabeled data containing both normal and abnormal samples can be utilized to enhance the model’s performance. Consequently, this paper focuses on addressing the challenging yet practical few-shot online anomaly detection and segmentation (FOADS) task. Under the FOADS framework, models are trained on a few-shot normal dataset, followed by inspection and improvement of their capabilities by leveraging unlabeled streaming data containing both normal and abnormal samples simultaneously.
To tackle this issue, we propose modeling the feature distribution of normal images using a Neural Gas network, which offers the flexibility to adapt the topology structure to identify outliers in the data flow. In order to achieve improved performance with limited training samples, we employ multi-scale feature embedding extracted from a CNN pre-trained on ImageNet to obtain a robust representation. Furthermore, we introduce an algorithm that can incrementally update parameters without the need to store previous samples. Comprehensive experimental results demonstrate that our method can achieve substantial performance under the FOADS setting, while ensuring that the time complexity remains within an acceptable range on MVTec AD and BTAD datasets.
@@
journal: Journal of XXXX
@@ 1 Introduction
Recent years have witnessed the development of anomaly detection and segmentation, particularly with the advent of deep learning. Early methods in the anomaly detection area primarily employ Auto-encoder [1, 2, 3, 4, 5] or GAN [6, 7, 8] to learn the normal distribution. Given a test image, these works try to reconstruct the image and compare the reconstructed images with the originals. Since the model is only trained on normal data, the anomaly parts are expected to be reconstructed poorly. Cohen et al.  [9] and Defard et al. [10] proposed the use of deep convolution neural network that trained on ImageNet dataset as feature extractor. Although the model is pre-trained on a general image classification dataset, it still offers strong performance on anomaly detection and segmentation even without any specific adaptation. In Table 1, we give a summary of part of the conventional methods.
@@
Figure 1: Illustration of the FOADS. The model $m_{0}$ is built on a training set which only contains $K$ normal images firstly. Subsequently, the model detects the test images $D^{1},D^{2},...,D^{n}$ and updates itself by leveraging the unlabeled streaming data which contains both normal and abnormal images. During the online learning process, our model assigns a pseudo label for each data and then trains itself with these labels.
However, there are still some unsolved limitations in these works. The core idea behind these methods is two-fold: first, feature matching between test images and normal images, and second, training a deep model to learn the normal distribution, both are data-driven. As a result, a substantial number of normal images are required to ensure training or matching accuracy. The effectiveness of these methods is significantly constrained when the training dataset contains only a small number of normal samples. However, in practical scenarios, in order to reduce interference with production processes, it is common that only limited high quality normal data is available for training an initial model.
@@
Another limitation is that most advanced methods ignore the the continuous data flow comprising a mixture of normal and abnormal data, generated in industrial scenarios after the model is deployed. If the model can autonomously and effectively utilize these unlabeled data to update its parameters, the practicality of the entire algorithm would be significantly enhanced. Consequently, there is a growing demand for models that require only limited samples for initial training and can enhance their detection and segmentation capabilities with the mixed data flow. We term this ability as few-shot online anomaly detection and segmentation (FOADS). The flowchart in Figure 1 depicts our proposed framework for detecting and updating by leveraging unlabelled data which contains both normal and abnormal instances.
@@
As mentioned above, existing works in anomaly detection and segmentation primarily leverage deep reconstruction models or memory banks. Although the reconstruction-based methods are very intuitive and interpretable, their performance is still constrained by the fact that Auto-encoder models often yield good reconstruction results for anomalous images as well [11]. Additionally, deep neural network models have been proven that they tend to overfit current data and lose knowledge learned before, which is called “catastrophic forgetting”  [12, 13]. Memory bank based approaches utilize cluster algorithm such as multivariate Gaussian model and coreset-subsampling to represent the entire feature space. Despite achieving higher accuracy than reconstruction-based methods, they are characterized by a predetermined isolated structure, making it difficult to adapt to unlabeled mixed data flow. Hence, the current anomaly detection and segmentation frameworks are not well-suited for directly addressing the FOADS problem. In the realm of online cluster learning, many researchers propose to leverage the Neural Gas (NG) [14] model, which can adjust its neighborhood size or relation according to changing scenes flexibly for better adaptation. But existing related work utilizing NG mainly concentrates on handcraft feature descriptors [15, 16, 17, 18] which are hard to design and less representative than multi-scale deep features whose performance has been demonstrated in previous studies [9, 10, 19]. Therefore, the intuitive idea is to introduce multi-scale deep features into previous NG networks to enhance the representational ability. However, the optimization algorithms of NG, such as Competitive Hebbian learning and Kohonen adjustment, are computationally intensive. The inherent limitation of these learning strategies, which update cluster centroid vectors one by one, results in an unacceptable running speed, severely limiting its practical value.
@@
MemAE [1] Proposing a memory-augmented autoencoder for anomaly detection Often yielding good reconstruction results for anomalous images and suffering from “catastrophic forgetting” in online learning
SMAI [5] Leveraging superpixel inpainting for better reconstruction.
@@
    We first recognize the importance of few-shot online anomaly detection and segmentation (FOADS) and define a problem setting to better organize the FOADS research. Compared with the widely studied anomaly detection and segmentation, FOADS is more challenging but more practical.
@@ 2.1 Anomaly Detection and Segmentation
Anomaly detection and segmentation are two important research topics that try to detect novel image and pixel after observing normal samples only. To tackle the challenge that no anomaly sample is available during training, a series of approaches have been proposed. Many existing works leveraging deep learning focus on generative models including Auto-encoder [1, 2, 3, 20] or Generative Adversarial Networks [6, 7, 8, 21]. The main idea of these approaches is training the networks on the normal data by minimizing the image reconstruction loss. The anomalous region can be detected by comparing the test image and the reconstructed one pixel by pixel since the model is only trained to rebuild normal images.
@@
However, the significant reconstruction error due to the diverse details in the image makes the accuracy considerably unsatisfactory. To address this issue, other researchers propose to model the normal features extracted by a CNN pre-trained on the ImageNet with a memory bank [9, 23] or probability density function [10, 24, 25]. Then the anomaly region can be detected by measuring the distance between test image and normal distribution. The Mahalanobis distance is often employed for measuring the distance between the test data and normal distribution, but the normal distribution does not always obey the Gaussian distribution which is the hypothesis of Mahalanobis distance. Therefore, some researchers suggest training a “Normalizing Flow” network to transform normal features into a standard normal distribution and give the anomaly score by deviation from standard mean [26, 27].
@@ 2.2 Online Learning and Detection
Online learning aims at handling streaming data, allowing the learning process to continue as data is collected [28, 29, 30]. A typical application of online learning is the detection in changing scenes, especially in video object detection. Kang et al. [31] propose a CNN based framework that leverages simple object tracking for video object detection. But this method consists of several different stages, resulting in high time complexity in video detection. Subsequently, a number of works use recurrent neural network (RNN) with long short-term memory (LSTM) cells [32] to replace the naive tractor [33, 34]. The feedback and gating mechanism of the LSTM block provide the model ability of processing relevant information under the temporal and spatial smoothness assumption [35].
In the realm of online anomaly detection, existing work primarily concentrates on video processing area [36, 15, 37, 38]. In order to enable the model to adjust its structure with the data flow, some previous work proposed an online Self-Organizing Map (online SOM) model and obtained some certain results. Nevertheless, the SOM has an inherent limitation that lattice structure is fixed which makes some null nodes far away from realistic samples. To address this limitation, some researchers propose a more flexible network called Neural Gas which learns data topology utilizing Hebbian learning.
@@ 2.3 Few-shot learning
Few-shot learning aims to enable the model to recognize unseen novel patterns while just leveraging very few training samples. To achieve few-shot learning, researchers usually adopt techniques like metric learning or optimization-based algorithm [39, 40, 41, 42, 43]. Optimization-based methods often employ a meta-learner which can generate weights for a new learner to process new tasks. In order to generate the meta-learner, researchers generally apply a separate network with a memory bank [44, 45] to produce learners or model it as an optimization procedure [46, 47]. Methods based on metric learning attempt to learn a feature embedding space where the same class is close while different classes are far apart. To build the distinguishable embedding space, Koch et al. [48] uses a Siamese network [49] which can minimize the distance between the same class and maximize otherwise. Then prototypical networks [42] that compute per-class representation and relation networks [41] which learn both an embedding and a distance function are proposed.
Some researchers extend few-shot learning to object detection and instance segmentation. Kang et al. [50] generate a meta-learner based on YOLOv2 [51] directly while Wang et al. propose a two stage method called TFA [52] which trains Faster R-CNN [53] and then only fine-tunes the predictor heads, achieving state-of-the-art results in object detection. In the aspect of the instance segmentation, Meta R-CNN [54] and Siamese Mask R-CNN [55] are developed to compute embedding of the train set and combine the embedding with feature map extracted by the backbone network.
@@
Figure 2: The pipeline diagram of the proposed algorithm. We first train the initial $K$-NG model on the few-shot training dataset. Then utilizing initial model for detecting online data and assigning pseudo labels to filter anomalous data. Finally, the remaining normal data is used to update parameters and adjust topology structures of the model.
@@ 3 Few-shot Online Anomaly Detection and Segmentation
Figure 3: The diagram illustrates the feature extraction pipeline. The images are sent into the pre-trained CNN to extract feature embeddings at different scales. The embeddings from different CNN blocks are concatenated and subsequently reduced in dimension through random selection.
@@ 3.1 Feature Extraction
Previous works utilizing Neural Gas for online clustering primarily focused on handcraft feature descriptors [15, 16, 17, 18]. In order to achieve competitive performance, we opt to utilize patch deep embeddings extracted by Convolution Neural Network (CNN) that has been pre-trained on the ImageNet dataset. Due to the presence of downsampling in CNN, each embedding in the CNN feature map is associated with a spatially corresponding patch of the image. An interpolation algorithm is then applied to obtain a high-resolution dense map
Feature maps extracted from different CNN layers contain information from different semantic levels and resolutions. To capture both fine-grained and global contexts, we concatenate feature embeddings from different layers. Similarly, as proposed by Defard et al. [10], we apply a simple random dimensionality reduction to decrease the complexity while maintaining high performance. The process schematic for the feature embedding extraction is depicted in Figure 3.
Figure 4: Comparison of two methods for characterizing the normal feature manifold.. Figure (a): The naive memory bank algorithm to model the normal manifold. Figure (b): The Neural Gas model leverages a graph to establish the topology relations among different feature descriptors. The topology structure endows the model a better capacity of representation.
@@ 3.2 Initialization of the Neural Gas
where the $\{x^{k}|k\in[1,N]\}$ are the feature embeddings assigned to the given neuron node, $N,\mu$ is the number of embeddings and sample mean respectively; the term $\epsilon I$ is the regularisation which makes the sample covariance matrix full rank and invertible. The threshold $T_{s_{i}}$ is calculated by using the mean distance between neuron $s_{i}$ and its neighboring neurons that are connected by edges:
$T_{s_{i}}=\underset{n\in\mathcal{N}_{s_{i}}}{\operatorname{mean}}\|\mathcal{A}_{s_{i}}-\mathcal{A}_{n}\|_{2},$ (3)
@@ 3.3 Online Learning of the Neural Gas
Figure 5: Diagram of the online learning process. Figure (a) shows the circumstance that the distance between $x$ and $s_{1}$ is greater than threshold $T_{s_{1}}$, in this case $x$ is abandoned. In Figure (b), the distance between $x$ and $s_{1}$ is less than threshold $T_{s_{1}}$ but there is no existing connection between $s_{1}$ and $s_{2}$, hence we add a new edge between them and assign $x$ to $s_{1}$. If there is existing connection between $s_{1}$ and $s_{2}$ like the situation in Figure (c), the age of the existing edge is set to 0.
@@ 3.4 Anomaly Detection Rule
This paper, assumes that the normal samples obey the Gaussian distribution and then leverages Mahalanobis distance to generate an anomaly score to each patch embedding. Defard et al. [10] first propose this distance measure in anomaly segmentation, but they restrict each patch embedding to the corresponding position in the image. When input images are not aligned strictly, modeling patch feature embeddings in fixed positions fails. In order to address this issue, we apply a global search in the NG network. Given a test image $I$ of size $W\times H$, we use $x_{ij}$ to denote the feature embedding of patch in position $(i,j)$. To calculate the anomaly score of position $(i,j)$, the nearest neuron in the network is searched firstly by:

### v2

@@ Abstract
Detecting anomaly patterns from images is a crucial artificial intelligence technique in industrial applications. Recent research in this domain has emphasized the necessity of a large volume of training data, overlooking the practical scenario where, post-deployment of the model, unlabeled data containing both normal and abnormal samples can be utilized to enhance the performance. Consequently, this paper focuses on addressing the challenging yet practical few-shot online anomaly detection and segmentation (FOADS) task. Under the FOADS framework, models are trained on a few-shot normal dataset, followed by inspection and improvement of their capabilities by leveraging unlabeled streaming data containing both normal and abnormal samples simultaneously. Since the data stream has no ground truth labels, the model needs to filter the anomalous data by detection to avoid contaminating its parameters, which presents a significant challenge when initial samples are insufficient. To tackle this issue, we propose modeling the feature distribution of normal images using a Neural Gas network, which offers the flexibility to adapt the topology structure to identify outliers in the data flow. In order to achieve improved performance with limited training samples, we employ multi-scale feature embedding extracted from a CNN pre-trained on ImageNet to obtain a robust representation. Furthermore, we introduce an algorithm that can incrementally update parameters without the need to store previous samples. Comprehensive experimental results demonstrate that our method can achieve substantial performance under the FOADS setting, while ensuring that the time complexity remains within an acceptable range on MVTec AD and BTAD datasets. Code is available at https://github.com/Whishing/K-NG
@@
journal: Journal of Knowledge-Based Systems
@@ 1 Introduction
Recent years have witnessed the development of anomaly detection and segmentation, particularly with the advent of deep learning. Early methods in the anomaly detection area primarily employ Auto-encoder [1, 2, 3, 4] or GAN [5, 6, 7] to learn the normal distribution. Given a test image, these works try to reconstruct the image and compare the reconstructed images with the originals. Since the model is only trained on normal data, the anomaly parts are expected to be reconstructed poorly. Cohen et al.  [8] and Defard et al. [9] proposed the use of deep convolution neural network that trained on ImageNet dataset as feature extractor. Although the model is pre-trained on a general image classification dataset, it still offers strong performance on anomaly detection and segmentation even without any specific adaptation. In Table 1, we give a summary of part of the conventional methods.
@@
Figure 1: Illustration of the FOADS. The model $m_{0}$ is built on a training set which only contains $K$ normal images firstly. Subsequently, the model detects the test images $D^{1},D^{2},...,D^{n}$ and updates itself by incorporating the unlabeled streaming data which contains both normal and abnormal images. Throughout the online learning process, our model assigns a pseudo label for each data and then trains itself with these labels.
@@
However, there are still some unsolved limitations in these works. Advanced methods [7, 9, 10, 11] for unsupervised industrial image anomaly detection and segmentation ignore the continuous data flow comprising a mixture of normal and abnormal data, generated in industrial scenarios after the model is deployed. Another limitation is that most current approaches need to collect a substantial number of normal images for training. From an industrial perspective, there is a reluctance to allocate significant time to collecting extensive training data. Thus, if the model can be established rapidly with limited initial training samples and constantly boosts its abilities to achieve better detection results, the practicality of the entire algorithm would be significantly enhanced. We term this ability as few-shot online anomaly detection and segmentation (FOADS). The flowchart in Figure 1 depicts our proposed framework for detecting and updating by leveraging unlabelled data which contains both normal and abnormal instances.
@@
However, due to the initial training samples are insufficient, the model needs to be modified frequently and significantly during the online learning. As mentioned above, existing works in anomaly detection and segmentation primarily leverage deep reconstruction models or memory banks. Deep neural network models have been proven to tend to overfit current data and lose knowledge learned before, which is called “catastrophic forgetting”  [12]. Memory bank-based approaches utilize cluster algorithms such as multivariate Gaussian model and coreset-subsampling to represent the entire feature space. They are characterized by a predetermined isolated structure, making it difficult to adapt to unlabeled mixed data flow. Hence, the current anomaly detection and segmentation frameworks are not well-suited for directly addressing the FOADS problem. In the realm of online cluster learning, many researchers propose to leverage the Neural Gas (NG) [13] model, which can adjust its neighborhood size or relation according to changing scenes flexibly for better adaptation. Consequently, we choose NG as the base model for addressing FOADS problem. But existing related work utilizing NG mainly concentrates on handcraft feature descriptors [14] which are hard to design and less representative than multi-scale deep features whose performance has been demonstrated in previous studies [8, 9, 10]. Therefore, the intuitive idea is to introduce multi-scale deep features into previous NG networks to enhance the representational ability. However, the optimization algorithms of NG, such as Competitive Hebbian learning and Kohonen adjustment, are computationally intensive. The inherent limitation of these learning strategies, which update cluster centroid vectors one by one, results in an unacceptable running speed, severely limiting its practical value.
@@
DiffusionAD [11] Leveraging diffusion model for better reconstruction. Often yielding good reconstruction results for anomalous images and suffering from “catastrophic forgetting” in online learning
@@
    We recognize the importance of few-shot online anomaly detection and segmentation (FOADS) and define a problem setting to better organize the FOADS research. Compared with the widely studied anomaly detection and segmentation, FOADS is more challenging but more practical.
@@ 2.1 Anomaly Detection and Segmentation
Anomaly detection and segmentation are two important research topics that try to detect novel image and pixel after observing normal samples only. To tackle the challenge that no anomaly sample is available during training, a series of approaches have been proposed. Many existing works leveraging deep learning focus on generative models including Auto-encoder [1, 2, 3, 4, 11] or Generative Adversarial Networks [5, 6, 7]. The main idea of these approaches is training the networks on the normal data by minimizing the image reconstruction loss. The anomalous region can be detected by comparing the test image and the reconstructed one pixel by pixel since the model is only trained to rebuild normal images.
@@
However, the significant reconstruction error due to the diverse details in the image makes the accuracy considerably unsatisfactory. To address this issue, other researchers propose to model the normal features extracted by a CNN pre-trained on the ImageNet with a memory bank [8, 10] or probability density function [9, 16]. Then the anomaly region can be detected by measuring the distance between test image and normal distribution. The Mahalanobis distance is often employed for measuring the distance between the test data and normal distribution, but the normal distribution does not always obey the Gaussian distribution which is the hypothesis of Mahalanobis distance. Therefore, some researchers suggest training a “Normalizing Flow” network to transform normal features into a standard normal distribution and give the anomaly score by deviation from standard mean [17, 18].
@@ 2.2 Online Learning and Detection
Online learning aims at handling streaming data, allowing the learning process to continue as data is collected [19, 20, 21]. A typical application of online learning is the detection in changing scenes, especially in video object detection. Kang et al. [22] propose a CNN based framework that leverages simple object tracking for video object detection. But this method consists of several different stages, resulting in high time complexity in video detection. Recently, Wu et al. [23] propose a plug-and-play module for label-efficient framework for online object detection from streaming video.
In the realm of online anomaly detection, existing work primarily concentrates on video processing area [24, 25]. In order to enable the model to adjust its structure with the data flow, some previous work proposed an online Self-Organizing Map (SOM) model and obtained some certain results. Nevertheless, the SOM has an inherent limitation that lattice structure is fixed which makes some null nodes far away from realistic samples. To address this limitation, some researchers propose a more flexible network called Neural Gas which learns data topology utilizing Hebbian learning.
@@ 2.3 Few-shot Learning
Few-shot learning aims to enable the model to recognize unseen novel patterns while just leveraging very few training samples. To achieve few-shot learning, researchers usually adopt techniques like metric learning or optimization-based algorithm [26, 27, 28]. Optimization-based methods often employ a meta-learner which can generate weights for a new learner to process new tasks. In order to generate the meta-learner, researchers generally apply a separate network [29] to produce learners or model it as an optimization procedure [30]. Methods based on metric learning attempt to learn a feature embedding space where the same class is close while different classes are far apart. To build the distinguishable embedding space, Shao et al. [31] uses a Siamese network which can minimize the distance between the same class and maximize otherwise. Then prototypical networks that compute per-class representation and relation networks [32] which learn both an embedding and a distance function are proposed.
Some researchers extend few-shot learning to object detection and instance segmentation. Kang et al. [33] generate a meta-learner based on YOLOv2 directly while Wang et al. propose a two stage method called TFA [34] which trains Faster R-CNN and then only fine-tunes the predictor heads, achieving state-of-the-art results in object detection. In the aspect of the instance segmentation, base-metal learner [35] and Siamese DETR [36] are developed to compute embedding of the train set and combine the embedding with feature map extracted by the backbone network.
@@
Figure 2: The pipeline diagram of the proposed algorithm. Firstly, the initial $K$-NG model is trained on the few-shot training dataset. Subsequently, the initial model is employed to detect online data and assign pseudo labels to filter anomalous data. Finally, the remaining normal data is used to update parameters and adjust topology structures of the model.
2.4 Few-shot Online Anomaly Detection
While there is a dearth of research on few-shot online anomaly detection and segmentation for industrial images, several existing literature has focused on detection of other data structures  [24, 37, 38]. Meng et al.  [37] propose leveraging a meta-knowledge learner for continuously adapting the model to new wave signal sources. Li et al.  [38] exploit a dueling triplet network to learn domain-invariant feature representation of time series and incremental update the classifier. However, these approaches are limited to predicting whether the sample is anomalous or not, which is not suitable for industrial image anomaly detection. In this paper, we not only provide the classification label but also predict per-pixel scores for accurate locating.
@@ 3 Few-shot Online Anomaly Detection and Segmentation
Figure 3: The diagram illustrates the feature extraction pipeline, wherein the images are input into a pre-trained CNN to extract feature embeddings at various scales. The embeddings obtained from different CNN blocks are concatenated and then dimensionally reduced through random selection.
@@ 3.1 Feature Extraction
Previous works utilizing Neural Gas for online clustering primarily focused on handcraft feature descriptors [14, 39]. In order to achieve competitive performance, we opt to utilize patch deep embeddings extracted by Convolution Neural Network (CNN) that has been pre-trained on the ImageNet dataset. Due to the presence of downsampling in CNN, each embedding in the CNN feature map is associated with a spatially corresponding patch of the image. An interpolation algorithm is then applied to obtain a high-resolution dense map
Feature maps extracted from different CNN layers contain information from different semantic levels and resolutions. To capture both fine-grained and global contexts, we concatenate feature embeddings from different layers. Specifically, as shown in the Table 2, features at the end of the first residual block conv2_x, the second residual block conv3_x, and the third residual block conv4_x from the ResNet18 backbone are adopted. Similarly, as proposed by Defard et al. [9], we apply a simple random dimensionality reduction to decrease the complexity while maintaining high performance. The process schematic for the feature embedding extraction is depicted in Figure 3.
Table 2: Structure of the ResNet18 based feature embedding extractor.
Layer Name Output Size Convolution Kernel
conv1 112 $\times$ 112 $\times$ 64 7 $\times$ 7, 64, stride 2
conv2_x 56 $\times$ 56 $\times$ 64 3 $\times$ 3 max pool, stride 2 $\left[\begin{array}[]{cc}3\times 3,&64\\ 3\times 3,&64\end{array}\right]\times 2$
conv3_x 28 $\times$ 28 $\times$128 $\left[\begin{array}[]{cc}3\times 3,&128\\ 3\times 3,&128\end{array}\right]\times 2$
conv4_x 14 $\times$ 14 $\times$ 256 $\left[\begin{array}[]{cc}3\times 3,&256\\ 3\times 3,&256\end{array}\right]\times 2$
conv5_x 7 $\times$ 7 $\times$ 512 $\left[\begin{array}[]{cc}3\times 3,&512\\ 3\times 3,&512\end{array}\right]\times 2$
fc 1 $\times$ 1 average pool, 1000-d fc, softmax
Figure 4: Comparison of two methods for characterizing the normal feature manifold: Figure (a) depicts the naive memory bank algorithm used to model the normal manifold, while Figure (b) illustrates the Neural Gas model, which utilizes a graph to establish topology relations among different feature descriptors. The topology structure of the Neural Gas model provides it with enhanced representational capacity.
@@ 3.2 Initialization of the Neural Gas
where the $\{x^{k}|k\in[1,N]\}$ are the feature embeddings assigned to the given neuron node, $N,\mu$ is the number of embeddings and sample mean respectively; the term $\epsilon I$ is the regularisation which makes the sample covariance matrix full rank and invertible. The threshold $T_{s_{i}}$ is calculated by using the median distance between neuron $s_{i}$ and its neighboring neurons that are connected by edges:
$T_{s_{i}}=\underset{n\in\mathcal{N}_{s_{i}}}{\operatorname{median}}\|\mathcal{A}_{s_{i}}-\mathcal{A}_{n}\|_{2},$ (3)
@@ 3.3 Online Learning of the Neural Gas
Figure 5: Diagram of the online learning process. Figure (a) shows the circumstance that the distance between $x$ and $s_{1}$ exceeds the threshold $T_{s_{1}}$, resulting in the abandonment of $x$. In Figure (b), the distance between $x$ and $s_{1}$ is less than threshold $T_{s_{1}}$, but there is no existing connection between $s_{1}$ and $s_{2}$. Therefore, a new edge is added between them, and $x$ is assigned to $s_{1}$. In the situation depicted in Figure (c), where there is an existing connection between $s_{1}$ and $s_{2}$, the age of the existing edge is reset to 0
@@ 3.4 Anomaly Detection Rule
[…]